# Symbolic Feedforward Networks for Probabilistic Finite Automata: Exact Simulation and Learnability

### v1

@@ 4.1 Encoding Input Strings into the Feedforward Network
We extend the symbolic input encoding mechanism proposed in Dhayalkar (2025d) to the probabilistic setting. Rather than embedding the entire input string $x=x_{1}x_{2}\cdots x_{L}\in\Sigma^{*}$ into a single vector, our construction treats each symbol as a control instruction that selects a specific stochastic transition operator to apply during the forward pass.
Transition Matrix Selection: For each symbol $x\in\Sigma$, we associate a probabilistic transition matrix $T^{x}\in\mathcal{S}_{n}$, where $T^{x}_{ij}=P(q_{j}\mid q_{i},x)$. These matrices can be:
@@
Feedforward Composition: Let $s_{t-1}\in\Delta^{n}\subseteq\mathbb{R}^{1\times n}$ be the state distribution after processing $x_{1}\cdots x_{t-1}$. For the current symbol $x_{t}$, the probabilistic update is:
$s_{t}=s_{t-1}T^{x_{t}},$
@@
Symbol-Driven Network Execution: The network processes each input symbol by selecting the corresponding $T^{x_{t}}$ and applying it to the current state distribution. Thus the mechanism is symbolic and dynamic.
Acceptance Computation: After processing the full string, the output vector $s_{L}\in\Delta^{n}$ reflects the final state distribution. The acceptance probability is computed as:
@@
Interpretation: This encoding generalizes the deterministic and nondeterministic constructions of Dhayalkar (2025d) by replacing binary state tracking with continuous probability mass propagation. It preserves the compositional, symbol-wise structure of automata while enabling soft, differentiable transitions that are compatible with standard feedforward networks.
Key Difference from NFAs: In contrast to the binary thresholded updates of Dhayalkar (2025d), our approach uses matrix-vector products over $\Delta^{n}$, resulting in a probabilistic forward pass that remains entirely symbolic, interpretable, and gradient-compatible.
Interpretation: This encoding generalizes the deterministic and nondeterministic constructions of Dhayalkar (2025d) by replacing binary state tracking with continuous probability mass propagation. It preserves the compositional, symbol-wise structure of automata while enabling soft, differentiable transitions that are compatible with standard feedforward networks.
Key Difference from NFAs: In contrast to the binary thresholded updates of Dhayalkar (2025d), our approach uses matrix-vector products over $\Delta^{n}$, resulting in a probabilistic forward pass that remains entirely symbolic, interpretable, and gradient-compatible.
@@ 4.2 Regular Language Recognition, Parameter Efficiency and Equivalence
The probabilistic feedforward architecture described above maintains a one-to-one correspondence with symbolic PFAs, and inherits many of the desirable theoretical properties of the NFA construction presented in Dhayalkar (2025d). In this subsection, we characterize the class of languages recognized by the model, and analyze its parameter efficiency.
@@
The feedforward network simulates a probabilistic finite automaton (PFA), which assigns a probability $P_{\mathcal{A}}(x)\in[0,1]$ to each string $x\in\Sigma^{*}$. This defines a stochastic language, i.e., a function $\Sigma^{*}\rightarrow[0,1]$. To turn this into a classical accept/reject language, one typically fixes a threshold $\tau\in(0,1)$ and defines:
@@
The resulting thresholded language $\mathcal{L}_{\tau}$ may or may not be regular, depending on the structure of the transition matrices and the choice of $\tau$. Unlike DFAs and NFAs—which always recognize regular languages—PFAs can induce non-regular languages under threshold semantics  Rabin and Scott (1959); Rabin (1963); Paz (1971). Our construction preserves this expressivity and defines a symbolic stochastic language through a feedforward architecture. This generalizes the NFA simulation framework of Dhayalkar (2025d), replacing binary reachability with smooth probabilistic computation.
@@
Let $\mathcal{A}=(Q,\Sigma,\delta,\pi_{0},F)$ be a probabilistic finite automaton with $|Q|=n$ states and $|\Sigma|=k$ symbols. Then any feedforward network that simulates $\mathcal{A}$ exactly via shared symbolic transition matrices requires only $O(kn^{2})$ parameters, independent of the input length $L$.
Proof. The complete proof is provided in Appendix A.5.
Explanation: This proposition guarantees that feedforward networks simulating PFAs scale only with the automaton size and alphabet, not with the input length. Unlike models that expand with string length or require dynamic memory, our construction uses shared symbolic operators for each input symbol, enabling input-length-independent parameterization.
The feedforward network maintains a fixed set of $k$ symbolic transition matrices—one per input symbol—and propagates state probabilities through matrix-vector multiplications. While the network depth scales with input length, the total parameter count remains constant. This parameter efficiency supports training and inference in resource-constrained settings, enables modular processing of new strings without architectural changes, and promotes strong generalization across sequence lengths, as the same symbolic operators apply recursively at each timestep.
Relation to Prior Work: Classical automata representations often require explicit unrolling or symbolic simulation per input Hopcroft et al. (2006); Rabin (1963), which limits scalability. Recurrent neural models, such as Neural Turing Machines Graves et al. (2014) or Adaptive Computation Time networks Graves (2017), frequently rely on increasing hidden state capacity or controller complexity to process longer sequences. In contrast, our method retains a fixed symbolic core that scales only with the alphabet size and state set, independent of input length, generalizing the parameter-sharing structure introduced in prior work on NFA simulation Dhayalkar (2025d).
@@
Let $L\subseteq\Sigma^{*}$ be any language recognized by a probabilistic finite automaton (PFA). Then:
@@
    (Forward Direction) For every PFA $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$ with $n$ states, there exists a symbolic feedforward network $f_{\theta}$ with:
@@
    such that for every string $x=x_{1}x_{2}\dots x_{L}\in\Sigma^{*}$, the network outputs $f_{\theta}(x)\in[0,1]$ and this value equals the probability that the PFA accepts $x$.
@@
    (Reverse Direction) Every such symbolic feedforward network $f_{\theta}$ composed of row-stochastic transition matrices and linear matrix compositions corresponds to a probabilistic finite automaton $\mathcal{A}^{\prime}$ such that for all $x\in\Sigma^{*}$,
@@
Proof. The complete proof is provided in Appendix A.6.
Explanation: This theorem establishes a formal equivalence between feedforward neural networks with normalized stochastic transitions and classical probabilistic finite automata. Just as nondeterminism in NFAs was simulated by binary thresholding of activations, stochastic transitions in PFAs are realized via matrix-vector multiplications with row-stochastic matrices—preserving probability mass at every step. The forward pass of the network corresponds exactly to computing a Markov chain distribution over states. Taken together with Proposition 5.1, this yields a complete expressivity-and-learnability characterization of PFAs within symbolic feedforward networks.
Relation to Prior Work: While previous work has demonstrated simulation of DFAs and NFAs using feedforward layers Dhayalkar (2025c; d), this result extends that framework to the probabilistic case.
@@ 5 Learnability of PFAs via Feedforward Networks
Let $\mathcal{A}=(Q,\Sigma,\delta,q_{0},F)$ be a probabilistic finite automaton (PFA) where $\delta:Q\times\Sigma\to\mathcal{D}(Q)$ is a stochastic transition function, and let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}$ be a dataset of input strings $x_{i}\in\Sigma^{*}$ labeled by their acceptance probability $y_{i}=P_{\mathcal{A}}(x_{i})$ under $\mathcal{A}$. Then, there exists a parameterized feedforward network $f_{\theta}$ constructed according to Theorem 4.10, in which the transition matrices $\{T^{x}\}_{x\in\Sigma}$ are initialized randomly and trained via gradient-based optimization over $\mathcal{D}$, such that the trained network $f_{\theta^{*}}$ can approximate the true acceptance probabilities of $\mathcal{A}$ to arbitrarily small empirical error on $\mathcal{D}$.
Proof. The complete proof is provided in Appendix A.7.
Explanation: While Theorem 4.10 establishes that any PFA can be simulated exactly by a symbolic feedforward network using row-stochastic transition matrices and linear updates, this proposition shows that such simulators can also be learned from data. That is, the underlying stochastic transition matrices of the PFA do not need to be known in advance—they can be discovered by optimizing a feedforward network against probabilistic acceptance labels.
This result shows that symbolic feedforward networks are not only capable of simulating known PFAs, but can also learn their behavior via supervised training. Given labeled examples of input strings and acceptance probabilities—potentially from an unknown PFA—the network can learn the underlying probabilistic transitions and decision boundaries using gradient-based optimization. This extends learnability guarantees from deterministic and nondeterministic automata Dhayalkar (2025d) to the probabilistic setting, demonstrating that regular stochastic languages can be represented and learned within standard neural architectures. Importantly, the parameter efficiency established in Proposition 4.9 still holds: the number of trainable parameters remains constant with respect to input length.
Relation to Prior Work: Unlike prior neural-symbolic models that approximate automata behavior heuristically or extract automata post hoc, our construction yields symbolic, interpretable, and trainable simulators. In contrast to black-box approximators or RNNs that lack interpretability and structural guarantees, our method maintains exact structure while enabling learnability. This continues the line of work in Dhayalkar (2025d), now extended to the probabilistic domain.
@@ 6.1 Experimental Setup
For each validation experiment, we compute binary acceptance decisions by thresholding the output score $\hat{y}$ against 0.5, and compare it to the ground-truth acceptance decision from the true PFA. We report the mean accuracy, standard deviation, and 95% confidence intervals using the Student’s $t$-distribution across 5 random seeds for each configuration. All experiments are implemented in PyTorch Paszke et al. (2019) and executed on an NVIDIA GeForce RTX 4060 GPU with CUDA acceleration.
Table 1: Validation results for symbolic simulation experiments across both configurations. Accuracy and confidence intervals computed over 5 random seeds using Student’s $t$-distribution.
@@
6.2: Proposition 4.1 and Remark 4.2 1 1.0000 0.0000 (1.0000, 1.0000)
6.2: Proposition 4.1 and Remark 4.2 2 1.0000 0.0000 (1.0000, 1.0000)
@@
First, we test whether the probabilistic state vector remains a valid distribution after each symbolic update. Specifically, we verify whether each intermediate state vector lies in the unit simplex: all entries are in $[0,1]$ and the vector sums to $1$. For each of 5 random seeds in both configurations, we instantiate a fresh PFA and evaluate the randomly generated input strings. Across all seeds and examples in both configurations, the state vector remained exactly valid as shown in Table 1. This confirms that symbolic updates via stochastic matrices preserve probabilistic semantics.
Second, we validate the semantic correctness of the state vector by checking whether it accurately encodes the marginal probability of being in each state after consuming an input string. For this, we compute the true marginal distribution via dynamic programming using the standard forward algorithm from HMMs. This avoids path enumeration and ensures exact probability tracking over reachable state sequences. We compare the resulting distribution to the symbolic state vector computed via matrix products. As shown in Table 1, across both configurations and all seeds, the two matched with perfect accuracy in every case. Together, these results confirm that the symbolic probabilistic state vector is both structurally and semantically correct, validating Proposition 4.1 and Remark 4.2.
@@
To validate Theorem 4.5, we test whether symbolic matrix-based updates match the full probabilistic evolution of the state vector at each timestep. For each of 5 random seeds in both configurations, we generate a fresh PFA and evaluate the randomly sampled input strings. At each timestep, we compare the symbolic simulator’s state vector to the ground-truth probabilistic trace computed directly from the PFA. As shown in Table 1, across all seeds and examples in both configurations, the simulator matched the probabilistic trace exactly at every timestep, yielding perfect accuracy. This confirms the correctness of the probabilistic subset construction under symbolic matrix composition.
@@
To validate Lemma 4.6, we simulate $\varepsilon$-closure in a PFA with only $\varepsilon$-transitions and compare the resulting marginal distribution to the fixed-point convergence of repeated matrix-vector updates. Instead of 5 random seeds as mentioned in Experimental Setup 6.1, we run it across 500 seeds. For each random seed, we sample a random stochastic $T^{\varepsilon}$ and a start state $q_{0}$, then iterate symbolic updates until convergence. As shown in Table 1, across both configurations and all seeds, the symbolic result matched the true marginal distribution exactly, confirming that the matrix iteration correctly computes the probabilistic $\varepsilon$-closure distribution.
@@
To validate Proposition 4.7, we evaluate whether a symbolic softmax-based feedforward simulator reproduces the accept/reject decisions of a ground-truth PFA. For each of 5 random seeds in both configurations, we generate a fresh PFA and the input strings, then compare simulator outputs to thresholded acceptance probabilities from the PFA (using a threshold of 0.5). As shown in Table 1, across both configurations, all seeds and examples, the simulator matched the PFA’s acceptance behavior exactly with perfect accuracy. This confirms that the symbolic network fully simulates the probabilistic automaton.
@@
To validate Theorem 4.10, we test whether a symbolic feedforward simulator reproduces the accept/reject decisions of a ground-truth PFA across randomly sampled strings. For each of 5 random seeds in both configurations, we instantiate a new PFA, construct a corresponding simulator, and compare outputs on 100 held-out test strings. As shown in Table 1, the simulator matched the PFA’s decisions (based on a threshold of 0.5) exactly across all examples and seeds in both configurations. This confirms functional equivalence between the feedforward network and its corresponding probabilistic automaton.
@@
To validate our central claim Proposition 5.1, we train a symbolic softmax-based feedforward model to approximate the accept/reject behavior of a ground-truth PFA. For each of 5 random seeds in both configurations, we instantiate a new PFA with the predefined number of states and randomly sampled stochastic transitions. We then generate the predefined number of training strings along with a held-out 100 test strings per seed, each consisting of symbol sequences as mentioned in the Experimental Setup 6.1. Each string is labeled as accepted or rejected by the PFA based on whether its acceptance probability exceeds a threshold of 0.5.
The model in both configurations is trained for 5 epochs using the Adam optimizer Kingma and Ba (2014) with learning rate $0.01$ and binary cross-entropy loss. During training, we apply a row-wise softmax normalization to the transition matrices to ensure they remain row-stochastic, in addition to a sigmoid activation to the final output layer. After training, test accuracy is computed by thresholding the model’s output and comparing against the PFA’s decision. As shown in Table 1, across all seeds, the learned model matched the ground-truth PFA’s decisions exactly in the first configuration, and achieved near perfect accuracy in the second configuration. This confirms that PFAs are exactly learnable by symbolic feedforward networks. Figure 1 shows that both training and test loss decrease consistently across all seeds in both configurations, confirming stable convergence and effective learning.

### v2

@@ 4.1 Encoding Input Strings into the Feedforward Network
We extend the symbolic input encoding mechanism proposed in Dhayalkar (2025b) to the probabilistic setting with $\varepsilon$-transitions. Rather than embedding the entire input string $x=x_{1}x_{2}\cdots x_{L}\in\Sigma^{*}$ into a single vector, our construction treats each symbol as a control instruction that selects a specific stochastic transition operator to apply during the forward pass, while also applying a finite $\varepsilon$-closure after each step.
Transition Matrix Selection: For each symbol $x\in\Sigma$, we associate a probabilistic transition matrix $T^{x}\in\mathcal{S}_{n}$, where $T^{x}_{ij}=P(q_{j}\mid q_{i},x)$. In addition, we define an $\varepsilon$-transition matrix $T^{\varepsilon}\in\mathcal{S}_{n}$ and $(T^{\varepsilon})^{*}\;=\;\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}$. These matrices can be:
@@
Feedforward Composition: Let $s_{t-1}\in\Delta^{n}\subseteq\mathbb{R}^{1\times n}$ be the state distribution after processing $x_{1}\cdots x_{t-1}$ and applying $\varepsilon$-closure. For the current symbol $x_{t}$, the probabilistic update is:
$s_{t}=s_{t-1}\,T^{x_{t}}(T^{\varepsilon})^{*},$
@@
Symbol-Driven Network Execution: The network processes each input symbol by selecting the corresponding $T^{x_{t}}$, multiplying it into the current state distribution, and applying the finite $\varepsilon$-closure $(T^{\varepsilon})^{*}$. Thus the mechanism is symbolic, dynamic, and exactly consistent with the semantics of PFAs.
Acceptance Computation: After processing the full string, the output vector $s_{L}\in\Delta^{n}$ reflects the final state distribution (after $\varepsilon$-closure). The acceptance probability is computed as:
@@
Interpretation: This encoding generalizes the nondeterministic constructions of Dhayalkar (2025b) by replacing binary state tracking with continuous probability mass propagation, while explicitly incorporating $\varepsilon$-transitions. It preserves the compositional, symbol-wise structure of automata while enabling soft, differentiable transitions that are compatible with standard feedforward networks.
Key Difference from NFAs: In contrast to the binary thresholded updates of Dhayalkar (2025b), our approach uses matrix-vector products over $\Delta^{n}$, resulting in a probabilistic forward pass that remains entirely symbolic, interpretable, and gradient-compatible.
@@ 4.2 Regular Language Recognition, Parameter Efficiency and Equivalence
The probabilistic feedforward architecture described above maintains a one-to-one correspondence with symbolic PFAs, and inherits many of the desirable theoretical properties of the NFA construction presented in Dhayalkar (2025b). In particular, since $\varepsilon$-closures are guaranteed to terminate in at most $n$ steps as mentioned in Lemma 4.6, the forward computation is always finite and efficiently realizable. In this subsection, we characterize the class of languages recognized by the model, and analyze its parameter efficiency.
@@
The feedforward network simulates a probabilistic finite automaton (PFA) with $\varepsilon$-transitions. The automaton assigns a probability $P_{\mathcal{A}}(x)\in[0,1]$ to each string $x\in\Sigma^{*}$. This defines a stochastic language, i.e., a function $\Sigma^{*}\rightarrow[0,1]$. To turn this into a classical accept/reject language, one typically fixes a threshold $\tau\in(0,1)$ and defines:
@@
The resulting thresholded language $\mathcal{L}_{\tau}$ may or may not be regular, depending on the structure of the transition matrices and the choice of $\tau$. Unlike DFAs and NFAs—which always recognize regular languages—PFAs can induce non-regular languages under threshold semantics (for some choices of transition probabilities and thresholds) Rabin and Scott (1959); Rabin (1963); Paz (1971). Our construction preserves this expressivity and defines a symbolic stochastic language through a feedforward architecture. This generalizes the NFA simulation framework of Dhayalkar (2025b), replacing binary reachability with smooth probabilistic computation.
@@
Let $\mathcal{A}=(Q,\Sigma\cup\{\varepsilon\},\delta,\pi_{0},F)$ be a probabilistic finite automaton with $|Q|=n$ states and $|\Sigma|=k$ symbols, where $\varepsilon$-closures terminate in at most $n$ steps. Then any feedforward network that simulates $\mathcal{A}$ exactly via shared symbolic transition matrices requires only $O(kn^{2})$ parameters, independent of the input length $L$.
Proof Sketch. The proof sketch is provided in Appendix A.5.
Explanation: This proposition guarantees that feedforward networks simulating PFAs scale only with the automaton size and alphabet, not with the input length. Unlike models that expand with string length or require dynamic memory, our construction uses shared symbolic operators for each input symbol, together with a single $\varepsilon$-transition operator applied through a finite closure of at most $n$ steps, enabling input-length-independent parameterization.
The feedforward network thus maintains a fixed set of $k$ symbolic transition matrices—one per input symbol—plus a single $\varepsilon$-transition matrix. The closure operation $(T^{\varepsilon})^{*}=\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}$ does not add parameters, since it is computed symbolically by repeated application of $T^{\varepsilon}$. While the network depth scales with input length, the total parameter count remains constant. This parameter efficiency supports training and inference in resource-constrained settings, enables modular processing of new strings without architectural changes, and promotes strong generalization across sequence lengths.
Relation to Prior Work: Classical automata representations often require explicit unrolling or symbolic simulation per input Hopcroft et al. (2006); Rabin (1963), which limits scalability. Recurrent neural models, such as Neural Turing Machines Graves et al. (2014) or Adaptive Computation Time networks Graves (2017), rely on increasing hidden state capacity or controller complexity to process longer sequences. In contrast, our method retains a fixed symbolic core that scales only with the alphabet size and state set, independent of input length, generalizing the parameter-sharing structure introduced in prior work on NFA simulation Dhayalkar (2025b).
@@
Let $\mathcal{L}\subseteq\Sigma^{*}$ be any language recognized by a probabilistic finite automaton (PFA) with $\varepsilon$-transitions, where $\varepsilon$-closures terminate in at most $n$ steps as mentioned in Lemma 4.6. Then:
@@
    (Forward Direction) For every PFA $\mathcal{A}=(Q,\Sigma\cup\{\varepsilon\},\delta,q_{0},F)$ with $n$ states, there exists a symbolic feedforward network $f_{\theta}$ with:
@@
    • a single $\varepsilon$-transition matrix $T^{\varepsilon}\in[0,1]^{n\times n}$,
@@
    such that for every string $x=x_{1}x_{2}\dots x_{L}\in\Sigma^{*}$, the network outputs $f_{\theta}(x)\in[0,1]$ and this value equals the probability that the PFA accepts $x$, where $\varepsilon$-closure is computed by $(T^{\varepsilon})^{*}=\sum_{m=0}^{n-1}(T^{\varepsilon})^{m}$.
@@
    (Reverse Direction) Every such symbolic feedforward network $f_{\theta}$ composed of row-stochastic transition matrices, a finite $\varepsilon$-closure operator, and linear matrix compositions corresponds to a probabilistic finite automaton $\mathcal{A}^{\prime}$ such that for all $x\in\Sigma^{*}$,
@@
Proof Sketch. The proof sketch is provided in Appendix A.6.
Explanation: This theorem establishes a formal equivalence between feedforward neural networks with normalized stochastic transitions (including a finite $\varepsilon$-closure) and classical probabilistic finite automata. Just as nondeterminism in NFAs was simulated by binary thresholding of activations Dhayalkar (2025b), stochastic transitions in PFAs are realized via matrix-vector multiplications with row-stochastic matrices, interleaved with finite $\varepsilon$-closure operations. The forward pass of the network corresponds exactly to computing a Markov chain distribution over states. Taken together with Proposition 5.1, this yields a complete expressivity-and-learnability characterization of PFAs within symbolic feedforward networks.
Relation to Prior Work: While previous work has demonstrated simulation of DFAs and NFAs using feedforward layers Dhayalkar (2025a; b), this result extends that framework to the probabilistic case.
@@ 5 Learnability of PFAs via Feedforward Networks
Let $\mathcal{A}=(Q,\Sigma\cup\{\varepsilon\},\delta,q_{0},F)$ be a probabilistic finite automaton (PFA) with $\varepsilon$-transitions, where $\delta:Q\times(\Sigma\cup\{\varepsilon\})\to\mathcal{D}(Q)$ is a stochastic transition function and $\varepsilon$-closures terminate in at most $n$ steps as mentioned in Lemma 4.6. Let $\mathcal{D}=\{(x_{i},y_{i})\}_{i=1}^{m}$ be a dataset of input strings $x_{i}\in\Sigma^{*}$ labeled by their acceptance probability $y_{i}=P_{\mathcal{A}}(x_{i})$ under $\mathcal{A}$. Then, there exists a parameterized feedforward network $f_{\theta}$ constructed according to Theorem 4.10, in which the transition matrices $\{T^{x}\}_{x\in\Sigma}$ and the $\varepsilon$-transition matrix $T^{\varepsilon}$ are initialized randomly and trained via gradient-based optimization over $\mathcal{D}$, such that the trained network $f_{\theta^{*}}$ can approximate the true acceptance probabilities of $\mathcal{A}$ to arbitrarily small empirical error on $\mathcal{D}$.
Proof Sketch. The proof sketch is provided in Appendix A.7.
Explanation: While Theorem 4.10 establishes that any PFA can be simulated exactly by a symbolic feedforward network using row-stochastic transition matrices, a single $\varepsilon$-transition matrix, and finite $\varepsilon$-closures, this proposition shows that such simulators can also be learned from data. That is, the underlying stochastic transition matrices of the PFA do not need to be known in advance—they can be discovered by optimizing a feedforward network against probabilistic acceptance labels.
This result shows that symbolic feedforward networks are not only capable of simulating known PFAs, but can also learn their behavior via supervised training. Given labeled examples of input strings and acceptance probabilities—potentially from an unknown PFA—the network can learn the underlying probabilistic transitions and decision boundaries using gradient-based optimization. This extends learnability guarantees from nondeterministic automata Dhayalkar (2025b) to the probabilistic setting, demonstrating that regular stochastic languages can be represented and learned within standard neural architectures. Importantly, the parameter efficiency established in Proposition 4.9 still holds: the number of trainable parameters remains constant with respect to input length.
Relation to Prior Work: Unlike prior neural-symbolic models that approximate automata behavior heuristically or extract automata post hoc, our construction yields symbolic, interpretable, and trainable simulators. In contrast to black-box approximators or RNNs that lack interpretability and structural guarantees, our method maintains exact structure while enabling learnability. This continues the line of work in Dhayalkar (2025b), now extended to the probabilistic domain.
@@ 6.1 Experimental Setup
For each validation experiment, we compute binary acceptance decisions by thresholding the output score $\hat{y}$ against the threshold of $0.5$, and compare it to the ground-truth acceptance decision from the true PFA. For each seed and configuration, we evaluate on 100 randomly generated test strings. We report the mean accuracy, standard deviation, and 95% confidence intervals using the Student’s $t$-distribution across 5 random seeds for each configuration. All experiments are implemented in PyTorch Paszke et al. (2019) and executed on an NVIDIA GeForce RTX 4060 GPU with CUDA acceleration.
Table I: Validation results for PFA symbolic simulation experiments across both configurations. Accuracy and confidence intervals computed over 5 random seeds using Student’s $t$-distribution.
@@
6.2: Proposition 4.1 1 1.0000 0.0000 (1.0000, 1.0000)
6.2: Proposition 4.1 2 1.0000 0.0000 (1.0000, 1.0000)
6.2: Remark 4.2 1 1.0000 0.0000 (1.0000, 1.0000)
6.2: Remark 4.2 2 1.0000 0.0000 (1.0000, 1.0000)
@@
First, we test whether the probabilistic state vector remains a valid distribution after each symbolic update. Specifically, we verify whether each intermediate state vector lies in the unit simplex: all entries are in $[0,1]$ and the vector sums to $1$. For each of 5 random seeds in both configurations, we instantiate a fresh PFA and evaluate the randomly generated input strings. Across all seeds and examples in both configurations, the state vector remained exactly valid (across the 100 randomly generated test strings) as shown in Table I. This confirms that symbolic updates via stochastic matrices preserve probabilistic semantics.
Second, we validate the semantic correctness of the state vector by checking whether it accurately encodes the marginal probability of being in each state after consuming an input string. For this, we compute the true marginal distribution via dynamic programming using the standard forward algorithm from HMMs. This avoids path enumeration and ensures exact probability tracking over reachable state sequences. We compare the resulting distribution to the symbolic state vector computed via matrix products. As shown in Table I, across both configurations and all seeds, the two matched with perfect accuracy in every case across the 100 randomly generated test strings. Together, these results confirm that the symbolic probabilistic state vector is both structurally and semantically correct, validating Proposition 4.1 and Remark 4.2.
@@
To validate Theorem 4.5, we test whether symbolic matrix-based updates match the full probabilistic evolution of the state vector at each timestep. For each of 5 random seeds in both configurations, we generate a fresh PFA and evaluate the randomly sampled input strings. At each timestep, we compare the symbolic simulator’s state vector to the ground-truth probabilistic trace computed directly from the PFA. As shown in Table I, across all seeds and examples in both configurations, the simulator matched the probabilistic trace exactly at every timestep, yielding perfect accuracy across the 100 randomly generated test strings. This confirms the correctness of the probabilistic subset construction under symbolic matrix composition.
@@
To validate Lemma 4.6, we simulate $\varepsilon$-closure in a PFA with only $\varepsilon$-transitions and compare the resulting marginal distribution to the fixed-point convergence of repeated matrix-vector updates. Instead of 5 random seeds as mentioned in Experimental Setup 6.1, we run it across 500 random seeds. For each random seed, we sample a random stochastic $T^{\varepsilon}$ and a start state $q_{0}$, then iterate symbolic updates until convergence. As shown in Table I, across both configurations and all seeds, the symbolic result matched the true marginal distribution exactly across the 100 randomly generated test strings, confirming that the matrix iteration correctly computes the probabilistic $\varepsilon$-closure distribution.
@@
To validate Proposition 4.7, we evaluate whether a symbolic softmax-based feedforward simulator reproduces the accept/reject decisions of a ground-truth PFA augmented with $\varepsilon$-transitions. For each of 5 random seeds in both configurations, we generate a fresh PFA where $\varepsilon$-transitions are added independently with probability $0.3$ per state, alongside the standard symbol-labeled transitions. The simulator applies the $\varepsilon$-closure at the start and after every symbol, consistent with Lemma 4.6, and we compare its outputs to thresholded acceptance probabilities from the PFA (using a threshold of $0.5$). As shown in Table I, across both configurations, all seeds and examples, the simulator matched the PFA’s acceptance behavior exactly with perfect accuracy across all the 100 randomly generated test strings. This confirms that the symbolic network fully simulates the probabilistic automaton, including the stochastic effect of $\varepsilon$-transitions.
@@
To validate Theorem 4.10, we test whether a symbolic feedforward simulator reproduces the accept/reject decisions of a ground-truth PFA augmented with $\varepsilon$-transitions. For each of 5 random seeds in both configurations, we instantiate a new PFA, where $\varepsilon$-transitions are added independently with probability $0.3$ per state alongside the stochastic symbol-labeled transitions, and construct a corresponding simulator. The simulator applies $\varepsilon$-closure both initially and after each symbol, ensuring consistency with Lemma 4.6. The setup is identical to Proposition 4.7, but here the results are interpreted as establishing bidirectional equivalence. As shown in Table I, across both configurations, seeds, and examples, the simulator matched the PFA’s decisions (based on a threshold of $0.5$) exactly across all the 100 randomly generated test strings. This confirms functional equivalence between the feedforward network and its corresponding probabilistic automaton, including the presence of $\varepsilon$-transitions.
@@
To validate our central claim Proposition 5.1, we train a symbolic softmax-based feedforward model to approximate the accept/reject behavior of a ground-truth PFA augmented with $\varepsilon$-transitions. For each of 5 random seeds in both configurations, we instantiate a new PFA with the predefined number of states, randomly sampled stochastic symbol-labeled transitions, and $\varepsilon$-transitions added independently with probability $0.3$ per state. We then generate the predefined number of training strings along with a held-out set of 100 test strings per seed, each consisting of symbol sequences as described in the Experimental Setup 6.1. Each string is labeled as accepted or rejected by the PFA based on whether its acceptance probability (after applying $\varepsilon$-closure at each step) exceeds a threshold of $0.5$.
The model in both configurations is trained for 5 epochs using the Adam optimizer Kingma and Ba (2014) with learning rate $0.01$ and binary cross-entropy loss. During training, we apply a row-wise softmax normalization to the transition matrices to ensure that they remain row-stochastic, in addition to a sigmoid activation to the final output layer. After training, test accuracy is computed by thresholding the model’s output and comparing against the PFA’s decision. As shown in Table I, across all seeds, the learned model matched the ground-truth PFA’s decisions exactly in the first configuration, and achieved near perfect accuracy in the second configuration. This confirms that PFAs are exactly learnable by symbolic feedforward networks. Fig. 1 shows that both training and test loss decrease consistently across all seeds in both configurations, confirming stable convergence and effective learning.